# The Future Outcome Reasoning and Confidence Assessment Benchmark

### v2

@@ Abstract
Forecasting is an important task in many domains, such as technology and economics. However existing forecasting benchmarks largely lack comprehensive confidence assessment, focus on limited question types, and often consist of artificial questions that do not align with real-world human forecasting needs. To address these gaps, we introduce FOReCAst (Future Outcome Reasoning and Confidence Assessment), a benchmark that evaluates models’ ability to make predictions and their confidence in them. FOReCAst spans diverse forecasting scenarios involving Boolean questions, timeframe prediction, and quantity estimation, enabling a comprehensive evaluation of both prediction accuracy and confidence calibration for real-world applications.
The Future Outcome Reasoning and Confidence Assessment Benchmark
Zhangdie Yuan University of Cambridge zy317@cam.ac.uk                        Zifeng Ding University of Cambridge zd320@cam.ac.uk                        Andreas Vlachos University of Cambridge av308@cam.ac.uk
@@
1 Introduction
Recent advances in large language models (LLMs) have significantly improved their performance across a wide range of natural language processing (NLP) tasks. Alongside these developments, various benchmarks and datasets have been introduced to effectively assess the capabilities of LLMs, particularly in terms of knowledge and reasoning (Zellers et al., 2019; Guo et al., 2023). Fact-based benchmarks, like TruthfulQA (Lin et al., 2022) evaluate LLMs based on factual correctness, focusing on tasks like retrieving and verifying facts that are known.
Forecasting is a crucial yet challenging task across various domains, including technology, economics, and public policy. Unlike tasks that rely on retrieving and verifying existing knowledge, forecasting requires predicting plausible outcomes for future events, often under uncertainty and incomplete information. This makes forecasting particularly difficult, as models must infer trends, assess probabilities, and adapt to new information. Several datasets have been introduced to evaluate LLMs’ forecasting capabilities. ForecastQA (Jin et al., 2020) uses a multiple-choice format where models predict future outcomes, but it lacks confidence assessment. AutoCast (Zou et al., 2022) incorporates confidence intervals, however its confidence estimates are not designed for forecasting. Other datasets such as ExpTime (Yuan et al., 2024a) are artificially generated from structured data, focusing on explainable event forecasting based on temporal knowledge graphs.
All of these aforementioned benchmarks ignore a crucial aspect of forecasting: confidence evaluation. Confidence plays a central role in forecasting, as predictions about unresolved events inherently lack definitive correctness at the time of evaluation. Predictions made with absolute certainty are undesirable, even if they ultimately prove to be correct, because they fail to account for the uncertain nature of future events. Moreover, miscalibrated confidence can lead to poor decision-making: overconfident yet incorrect forecasts may result in costly errors, while underconfident but accurate predictions can erode trust in the model. Therefore, well-calibrated confidence scores are as crucial as the accuracy of the predictions themselves.
To address these gaps, we present FOReCAst: Future Outcome Reasoning and Confidence Assessment. FOReCAst focuses on three distinct types of forecasting questions, shown in Table 1: (1) Boolean questions, such as "Will there be a Frontier AI lab in China before 2026?"; (2) Timeframe Prediction, such as "When will OpenAI announce GPT-5?"; and (3) Quantity Estimation, such as "How many spacecrafts will land on the moon in each of the following years?" We conduct experiments using a range of models differing in size, training objectives, and cutoff times, and explore multiple methods for estimating model confidence. Our results reveal that forecasting remains highly challenging for current LLMs, particularly in confidence evaluation, with no direct correlation between prediction performance and confidence calibration, and while larger models sometimes improve performance, the effect is inconsistent.
@@ 2 FOReCAst: Problem Formulation
System responses in FOReCAst consist of (1) a prediction answering a question given the available information and (2) a confidence score in the prediction. This
ensures a comprehensive assessment of forecasting, accounting for both correctness and confidence calibration.
@@
$M(Q)\to(A,C),\quad C\in[0,1].$ (1)
$A=\arg\max_{a\in\mathcal{A}}P(X=a|Q,\mathcal{K}(t_{train})).$ (2)
@@ 3 Evaluating Predictions and Confidence
Confidence calibration is assessed using a modified version of the Brier score (Brier, 1950), which measures the mean squared error between predicted confidence and the gold confidence provided in the dataset, which we assume is provided, and
represents the likelihood of an event occurring. The modified Brier score is defined as:
@@
$\sigma^{\text{pred}}_{i}=\sigma_{\max}\cdot(1-C^{\text{pred}}_{i})+\sigma_{\min}\cdot C^{\text{pred}}_{i},$ (5)
$\sigma^{\text{gold}}_{i}=\sigma_{\max}\cdot(1-C^{\text{gold}}_{i})+\sigma_{\min}\cdot C^{\text{gold}}_{i}.$ (6)
@@
$E^{\text{APE}}_{i}=\frac{2}{1+e^{-\alpha\frac{|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}{Q^{\text{gold}}_{i}+\epsilon}}}-1,$ (8)
$E^{\text{MAE}}_{i}=\frac{2}{1+e^{-\alpha|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}}-1.$ (9)
@@ 4.1 Data Source and Question Selection
FOReCAst is constructed from Metaculus,111www.metaculus.com. Examples in Appendix A. an online forecasting platform where forecasters submit probabilistic predictions to questions across various domains. Metaculus aggregates individual forecasts into a continuously updated community prediction, which is finalized just before resolution. Each question has predefined resolution criteria, ensuring verifiable outcomes. To ensure dataset reliability, we include only questions with a definitive resolution and at least 100 forecasts to maintain statistical reliability. Ambiguous or subjectively resolved questions are excluded, and we remove those whose outcomes depend on arbitrary or uncontrollable factors. These steps ensure that FOReCAst consists of high-quality, well-formed forecasting questions with verifiable outcomes.
@@ 4.2 Extracting Confidence from Crowdsourced Forecasts
Formal Definition of Gold Confidence.
Gold confidence in FOReCAst is derived from the final Metaculus community prediction before resolution. Instead of directly using the predicted probability for the correct outcome, we compute a log score relative to a uniform baseline, ensuring that confidence reflects how much the forecast deviates from random guessing. This transformation prevents extreme probabilities in inherently uncertain scenarios and makes confidence scores more comparable across different forecasting tasks. The final score is mapped to $(0,1)$ using a sigmoid function.
@@
where $\sigma(x)$ is the sigmoid function.
Similarly, for timeframe prediction and quantity estimation, where the human-forecasted probability density function (PDF) assigns probability to a continuous outcome $x^{\text{gold}}$, gold confidence is computed as:
@@
where $f(x^{\text{gold}})$ is the forecasted probability density at the resolved outcome, and $f_{\text{uniform}}$ is the uniform baseline density over the valid range of values. The denominator 2 ensures numerical stability and scales confidence appropriately.
@@ 4.3 Dataset Statistics and Comparison
ForecastQA MCQ
AutoCast Various
ExpTime Boolean
@@ 5.1 Experimental Setup
GPT-2 GPT-2, GPT-2 XL 2017-12-01
Pythia 14M, 160M, 2.8B 2020-03-01
BLOOM 560M, 7B1 2021-12-01
LLaMA LLaMA-7B 2022-08-01
OLMo 1B, 7B, 7B-Instruct 2023-03-01
OLMo-2 7B, 7B-Instruct 2023-12-01
Table 3: Models used in our experiments, grouped by family and ordered by training data cutoff date, including: GPT-2 (Radford et al., 2019), Pythia (Biderman et al., 2023), BLOOM (Scao et al., 2023), LLaMA (Touvron et al., 2023), OLMo (Groeneveld et al., 2024), and OLMo-2 (OLMo et al., 2024).
@@ 5.2 Results and Findings
Our experiments on the FOReCAst dataset reveal that forecasting remains highly challenging for current LLMs, particularly in confidence estimation. While models achieve reasonable accuracy in point predictions, their uncertainty estimates—captured by Brier score and CRPS—vary significantly. This suggests that confidence evaluation must be treated as a separate challenge from prediction assessment.
Boolean Questions:
Table 4 shows that while some models achieve reasonable accuracy, their calibration remains inconsistent. For example, GPT2-XL achieves an accuracy of 0.67 but has a relatively high Brier score of 0.45, while OLMo-7B-Instruct achieves similar accuracy (0.65) with a lower Brier score of 0.38, suggesting better confidence estimation. Notably, models with later training cutoffs do not always outperform older ones; for instance, LLaMA-7B has lower accuracy (0.57) and a worse Brier score (0.52) than some earlier models. This indicates that forecasting accuracy depends on multiple factors beyond knowledge recency.
Model N ADE ($\downarrow$) CRPS ($\downarrow$)
GPT2 26 0.99 0.99
GPT2-XL 26 1.00 1.00
Pythia-14m 25 1.00 1.00
Pythia-160m 25 1.00 1.00
Pythia-2.8b 25 0.97 0.96
Bloom-560m 12 1.00 1.00
Bloom-7b1 12 1.00 1.00
Llama-7b 10 0.98 0.98
OLMo-1B 6 1.00 1.00
OLMo-7B 6 1.00 1.00
OLMo-7B-Instruct 6 1.00 1.00
OLMo-2-7B 4 1.00 1.00
OLMo-2-7B-Instruct 4 0.87 0.82
Table 5: Performance of forecasting models on Timeframe Prediction tasks in FOReCAst. Metrics include the number of evaluated questions (N), the normalized absolute days error (ADE), and the Continuous Ranked Probability Score (CRPS), where lower values indicate better performance.
Model N APE ($\downarrow$) MAE ($\downarrow$) CRPS ($\downarrow$)
GPT2 81 0.23 0.87 0.86
GPT2-XL 81 0.03 0.75 0.72
Pythia-14m 77 0.17 0.88 0.85
Pythia-160m 77 0.03 0.78 0.76
Pythia-2.8b 77 0.07 0.81 0.79
Bloom-560m 43 0.05 0.73 0.71
Bloom-7b1 43 0.06 0.75 0.72
Llama-7b 32 0.06 0.64 0.61
OLMo-1B 23 0.21 0.79 0.76
OLMo-7B 23 0.22 0.84 0.82
OLMo-7B-Instruct 23 0.24 0.80 0.77
OLMo-2-7B 20 0.02 0.65 0.62
OLMo-2-7B-Instruct 20 0.09 0.60 0.57
@@
Table 6: Performance of forecasting models on Quantity Estimation tasks in FOReCAst. Metrics include the number of evaluated questions (N), normalized absolute percentage error (APE), mean absolute error (MAE), and Continuous Ranked Probability Score (CRPS), where lower values are better.
Timeframe Prediction:
Table 5 reveals that most models struggle with predicting event timing, as both ADE and CRPS remain near the worst-case scenario of 1.0. Across all models, only one achieves a CRPS below 0.9: OLMo-2-7B-Instruct (CRPS = 0.82), which still indicates substantial uncertainty. Even models with more recent training data, such as OLMo-2-7B (CRPS = 0.99), fail to make well-calibrated temporal forecasts. These results suggest that even when models correctly anticipate whether an event will happen, quantifying when it will occur remains a major challenge.
Quantity Estimation:
Table 6 highlights that models differ widely in their ability to estimate numerical values. Some achieve relatively low APE, such as GPT2-XL (APE = 0.03) and Pythia-160M (APE = 0.02), yet their confidence calibration, CRPS, does not always align with their point prediction performance. For instance, GPT2-XL has a CRPS of 0.72, while Pythia-160M has a slightly higher CRPS of 0.76, despite achieving lower APE. In contrast, OLMo-2-7B-Instruct, which has a relatively higher APE of 0.08, achieves the lowest CRPS (0.57) among all models. These results indicate that prediction quality and confidence calibration do not necessarily improve together, reinforcing the complexity of numerical forecasting.
@@
Larger models do not consistently improve forecasting performance. Within the Pythia family, Pythia-2.8b shows occasional gains in accuracy and calibration over Pythia-14m and Pythia-160m, but the improvements are not uniform across all metrics. Similarly, while OLMo-2-7B variants sometimes achieve lower CRPS and MAE in quantity estimation, these benefits often come with trade-offs in point prediction. These results suggest that model size alone is not a reliable predictor of forecasting performance. This highlights the importance of developing task-specific techniques rather than relying on ever-larger models to solve the forecasting problem.
@@
Table 7 compares base and instruct-tuned variants of OLMo-7B and OLMo-2-7B across Boolean, timeframe, and quantity forecasting tasks. For Boolean questions, OLMo-7B-Instruct achieves higher accuracy (0.65 vs. 0.21) and a lower Brier score (0.38 vs. 0.82), indicating better confidence calibration. In timeframe prediction, the instruct-tuned OLMo-2-7B-Instruct improves uncertainty estimation, with an ADE of 0.87 and CRPS of 0.82, compared to 0.9981 and 0.9970 for the base OLMo-2-7B. For quantity estimation, instruct-tuned models have slightly higher APE but lower MAE and CRPS, suggesting better uncertainty calibration. These results indicate that instruction tuning enhances confidence estimation, even if it does not always improve point prediction accuracy. This suggests a trade-off where instruct-tuned models prioritize more reliable uncertainty quantification.
@@
Table 8 compares different aggregation methods for deriving the final prediction and confidence estimate from the top 10 outputs of Llama-7B. Bayesian Aggregation achieves the highest accuracy (0.5796) and F1 score (0.3485), suggesting it is the most effective at identifying correct outcomes. However, Weighted Average yields a significantly lower Brier score (0.2914), indicating superior confidence calibration compared to other methods.
Majority Vote, Highest Confidence, and Logit Mean Probability produce comparable accuracy and F1 scores but have noticeably higher Brier scores, suggesting weaker uncertainty estimation. These results highlight that even when point prediction performance is similar, aggregation methods substantially impact confidence reliability. The challenge remains in developing techniques that optimize both accuracy and calibration simultaneously, emphasizing the importance of uncertainty-aware forecasting.

### v3

# Introducing FOReCAst: The Future Outcome Reasoning and Confidence Assessment Benchmark

@@ Abstract
Forecasting is an important task in many domains, such as technology and economics. However existing forecasting benchmarks largely lack comprehensive confidence assessment, focus on limited question types, and often consist of artificial questions that do not align with real-world human forecasting needs. To address these gaps, we introduce FOReCAst (Future Outcome Reasoning and Confidence Assessment), [^1] Publicly availble at https://huggingface.co/datasets/MoyYuan/FOReCAst a benchmark that evaluates models’ ability to make predictions and their confidence in them. FOReCAst spans diverse forecasting scenarios involving Boolean questions, timeframe prediction, and quantity estimation, enabling a comprehensive evaluation of both prediction accuracy and confidence calibration for real-world applications.
1 Introduction
Recent advances in large language models (LLMs) have significantly improved their performance across a wide range of natural language processing (NLP) tasks. Alongside these developments, various benchmarks and datasets have been introduced to effectively assess the capabilities of LLMs, particularly in terms of knowledge and reasoning (Zellers et al., 2019; Guo et al., 2023). Fact-based benchmarks, such as TruthfulQA (Lin et al., 2022), evaluate LLMs based on factual correctness, focusing on tasks like retrieving and verifying facts that are already known.
Forecasting is a crucial yet challenging task across various domains, including technology, economics, and public policy. Unlike tasks that rely on retrieving and verifying existing knowledge, forecasting requires predicting plausible outcomes for future events, often under uncertainty and incomplete information. This makes forecasting particularly difficult, as models must infer trends, assess probabilities, and adapt to new information. Several datasets have been introduced to evaluate LLMs’ forecasting capabilities. ForecastQA (Jin et al., 2020) uses a multiple-choice format where models predict future outcomes, but it lacks confidence assessment. AutoCast (Zou et al., 2022) incorporates confidence intervals; however, its confidence estimates are not specifically designed for forecasting tasks. Other datasets, such as ExpTime (Yuan et al., 2024a), are artificially generated from structured data and focus on explainable event forecasting based on temporal knowledge graphs.
All of these aforementioned benchmarks overlook a crucial aspect of forecasting: confidence evaluation. Confidence plays a central role in forecasting, as predictions about unresolved events inherently lack definitive correctness at the time of evaluation. Predictions made with absolute certainty are undesirable, even if they ultimately prove to be correct, because they fail to account for the uncertain nature of future events. Moreover, miscalibrated confidence can lead to poor decision-making: overconfident yet incorrect forecasts may result in costly errors, while underconfident but accurate predictions can erode trust in the model. Therefore, well-calibrated confidence scores are as crucial as the accuracy of the predictions themselves.
To address these gaps, we present FOReCAst: Future Outcome Reasoning and Confidence Assessment. FOReCAst focuses on three distinct types of forecasting questions, shown in Table 1: (1) Boolean Questions, such as "Will there be a Frontier AI lab in China before 2026?"; (2) Timeframe Prediction, such as "When will OpenAI announce GPT-5?"; and (3) Quantity Estimation, such as "How many spacecrafts will land on the moon in each of the following years?" We conduct experiments using a range of models differing in size, training objectives, and training cutoff times, and explore multiple methods for estimating model confidence. Our results reveal that forecasting remains highly challenging for current LLMs, particularly in confidence evaluation, with no consistent correlation between prediction accuracy and calibration. While larger models sometimes improve performance, the effect is inconsistent and highly task-dependent.
@@ 2 FOReCAst: Problem Formulation
System responses in FOReCAst consist of (1) a prediction answering a question given the available information and (2) a confidence score in the prediction. This ensures a comprehensive assessment of forecasting, accounting for both correctness and confidence calibration.
@@
$M(Q)\to(A,C)\quad where\quad A=\arg\max_{a\in\mathcal{A}}P(X=a|Q,\mathcal{K}(t_{train}))\quad and\quad C\in[0,1].$ (1)
@@ 3 Evaluating Predictions and Confidence
Confidence calibration is assessed using a modified version of the Brier score (Brier, 1950), which measures the mean squared error between predicted confidence and the gold confidence provided in the dataset, which we assume is provided, and represents the likelihood of an event occurring. The modified Brier score is defined as:
@@
$\sigma^{\text{pred}}_{i}=\sigma_{\max}\cdot(1-C^{\text{pred}}_{i})+\sigma_{\min}\cdot C^{\text{pred}}_{i},\quad\sigma^{\text{gold}}_{i}=\sigma_{\max}\cdot(1-C^{\text{gold}}_{i})+\sigma_{\min}\cdot C^{\text{gold}}_{i}.$ (4)
@@
$E^{\text{APE}}_{i}=\frac{2}{1+e^{-\alpha\frac{|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}{Q^{\text{gold}}_{i}+\epsilon}}}-1,\quad E^{\text{MAE}}_{i}=\frac{2}{1+e^{-\alpha|Q^{\text{pred}}_{i}-Q^{\text{gold}}_{i}|}}-1.$ (6)
@@ 4.1 Data Source and Question Selection
FOReCAst is constructed from Metaculus,222www.metaculus.com. Examples in Appendix A. an online forecasting platform where forecasters submit probabilistic predictions to questions across various domains. Metaculus aggregates individual forecasts into a continuously updated community prediction, which is finalized just before resolution. Each question has predefined resolution criteria, ensuring verifiable outcomes. To ensure dataset reliability, we include only questions with a definitive resolution and at least 100 forecasts to maintain statistical reliability. Ambiguous or subjectively resolved questions are excluded, and we remove those whose outcomes depend on arbitrary or uncontrollable factors. These steps ensure that FOReCAst consists of high-quality, well-formed forecasting questions with verifiable outcomes. The dataset spans a diverse range of domains such as politics, economics, health, and technology, with questions distributed across a wide temporal range; detailed statistics are provided in Table 9.
@@ 4.2 Extracting Confidence from Crowdsourced Forecasts
Formal Definition of Gold Confidence
Gold confidence in FOReCAst is derived from the final Metaculus community prediction just before question resolution. Rather than directly using the predicted probability for the correct outcome, we compute a log score relative to a uniform baseline, following Metaculus’s established approach for evaluating probabilistic forecasts  [^3] https://www.metaculus.com/help/scores-faq/. This transformation ensures that confidence reflects not only correctness, but also the amount of informational gain over random guessing. It provides a consistent interpretation of confidence across tasks and scales, particularly when outcomes lie in continuous spaces.
@@
where $\sigma(x)$ is the sigmoid function. This yields a score of 0.5 for a random-guess forecast ($P^{\text{gold}}=0.5$), and smoothly increases or decreases with higher or lower confidence.
For timeframe prediction and quantity estimation, where the human-forecasted probability density function (PDF) assigns likelihood to a continuous outcome $x^{\text{gold}}$, gold confidence is computed as:
@@
where $f(x^{\text{gold}})$ is the forecasted probability density at the resolved outcome, and $f_{\text{uniform}}$ is the uniform baseline density over the valid outcome range. The denominator of 2 stabilizes the scale and avoids over-amplifying minor density differences. This approach accounts for the resolution granularity of each question and ensures confidence is not inflated in uncertain or diffuse distributions.
While this scoring formulation is tailored to Metaculus, it can be adapted or replaced if used with other forecasting platforms or scoring protocols. Our goal is to align confidence with meaningful deviations from uncertainty, while maintaining a consistent and interpretable signal across all tasks.
@@ 4.3 Dataset Statistics and Comparison
ForecastQA (Jin et al., 2020) MCQ
AutoCast (Zou et al., 2022) Various
ExpTime (Yuan et al., 2024a) Boolean
@@ 5.1 Experimental Setup
GPT-2 (Radford et al., 2019) GPT-2, GPT-2 XL 2017-12-01
Pythia (Biderman et al., 2023) 14M, 160M, 2.8B 2020-03-01
BLOOM (Scao et al., 2023) 560M, 7B1 2021-12-01
LLaMA (Touvron et al., 2023) LLaMA-7B 2022-08-01
OLMo (Groeneveld et al., 2024) 1B, 7B, 7B-Instruct 2023-03-01
OLMo-2 (OLMo et al., 2024) 7B, 7B-Instruct 2023-12-01
Table 3: Models used in our experiments, grouped by family and ordered by training data cutoff date.
@@ 5.2 Results, Findings, and Discussions
Our experiments on the FOReCAst dataset reveal that forecasting remains a challenging task for current LLMs, particularly in estimating and expressing uncertainty. While models often achieve reasonable accuracy in their point predictions, their confidence calibration—measured by Brier score and CRPS—shows substantial variability. This indicates that uncertainty estimation should be treated as a distinct evaluation axis, not inferred from accuracy alone. Throughout the following sections, we offer interpretations and possible explanations for the observed behaviors.
Boolean Questions
Table 4 shows that while some models reach reasonable accuracy on binary classification tasks, their confidence calibration remains inconsistent. For instance, GPT2-XL achieves a relatively strong accuracy of 0.67 but has a higher Brier score of 0.45, whereas OLMo-7B-Instruct achieves a similar accuracy (0.65) with a lower Brier score of 0.38, indicating better alignment between predicted probabilities and actual outcomes. This discrepancy suggests that models may differ not only in their ability to identify the correct answer, but also in how they distribute confidence over binary options.
Importantly, model performance does not appear to correlate straightforwardly with scale or training recency. For example, LLaMA-7B, a more recent model, shows lower accuracy (0.57) and higher Brier score (0.52) than older models like GPT2-XL. This supports the observation that accuracy in binary forecasting tasks depends on a combination of world knowledge, probabilistic reasoning, and how well models can handle epistemic uncertainty, particularly when questions are unresolved or partially underspecified at prompt time. The gap between classification accuracy and calibration further underscores the importance of evaluating both, as good classification alone may mask unreliable probability estimates.
Quantity Estimation Timeframe Prediction
Model N APE ($\downarrow$) MAE ($\downarrow$) CRPS ($\downarrow$) Model N ADE ($\downarrow$) CRPS ($\downarrow$)
GPT2 81 0.23 0.87 0.86 GPT2 26 0.99 0.99
GPT2-XL 81 0.03 0.75 0.72 GPT2-XL 26 1.00 1.00
Pythia-14m 77 0.17 0.88 0.85 Pythia-14m 25 1.00 1.00
Pythia-160m 77 0.03 0.78 0.76 Pythia-160m 25 1.00 1.00
Pythia-2.8b 77 0.07 0.81 0.79 Pythia-2.8b 25 0.97 0.96
Bloom-560m 43 0.05 0.73 0.71 Bloom-560m 12 1.00 1.00
Bloom-7b1 43 0.06 0.75 0.72 Bloom-7b1 12 1.00 1.00
Llama-7b 32 0.06 0.64 0.61 Llama-7b 10 0.98 0.98
OLMo-1B 23 0.21 0.79 0.76 OLMo-1B 6 1.00 1.00
OLMo-7B 23 0.22 0.84 0.82 OLMo-7B 6 1.00 1.00
OLMo-7B-Instruct 23 0.24 0.80 0.77 OLMo-7B-Instruct 6 1.00 1.00
OLMo-2-7B 20 0.02 0.65 0.62 OLMo-2-7B 4 1.00 1.00
OLMo-2-7B-Instruct 20 0.09 0.60 0.57 OLMo-2-7B-Instruct 4 0.87 0.82
@@
Table 5: Performance of forecasting models on Quantity Estimation (left) and Timeframe Prediction (right) tasks in FOReCAst. All metrics are lower-is-better. APE: Absolute Percentage Error, MAE: Mean Absolute Error, CRPS: Continuous Ranked Probability Score, ADE: Absolute Days Error.
Timeframe Prediction
Table 5 demonstrates that timeframe prediction—estimating when an event will occur—is especially difficult for current models. Across the board, both ADE and CRPS values remain close to 1.0, indicating little improvement over naive or uninformed baselines. Only OLMo-2-7B-Instruct achieves a CRPS noticeably below this range (0.82), but even this is far from well-calibrated. The generally poor performance suggests that temporal forecasting presents challenges beyond simple factual retrieval or classification.
This difficulty likely reflects the inherent complexity of temporal prediction, which often requires models to reason over event dynamics, causality, and elapsed time—all aspects that are underrepresented in standard pretraining corpora. In contrast to Boolean tasks, where resolution criteria are clearly defined and outcomes are binary, timeframe questions require mapping language input to a continuous temporal target, which demands a finer-grained understanding of how events evolve over time. That most models fail to depart meaningfully from baseline-level CRPS indicates that LLMs may lack sufficient temporal inductive biases or exposure to structured time-referenced reasoning.
Quantity Estimation
As shown in Table 5, quantity estimation—predicting a numerical outcome—exhibits a broader range of performance across models. Some models such as Pythia-160M and GPT2-XL achieve very low absolute percentage error (APE = 0.02 and 0.03 respectively), indicating strong point prediction capability. However, their CRPS values (0.76 and 0.72) suggest relatively poor calibration. In contrast, OLMo-2-7B-Instruct, with a higher APE of 0.08, attains the lowest CRPS (0.57), suggesting it better captures uncertainty even if its point estimates are less precise.
This divergence illustrates that point prediction accuracy and confidence calibration are not necessarily aligned. Predicting a correct numerical value does not guarantee that a model’s probability distribution around its prediction reflects uncertainty appropriately. Since CRPS penalizes both distance from the true value and over/under-confidence, this metric provides a more comprehensive view of performance in numerical forecasting tasks. These results suggest that even relatively strong predictive models may require additional mechanisms—such as structured uncertainty modeling or calibration-sensitive objectives—to produce well-calibrated forecasts in quantity estimation settings.
@@
Model size alone does not reliably predict forecasting performance across tasks. Within the Pythia family, Pythia-2.8b occasionally improves over smaller variants in certain metrics—for example, it achieves lower CRPS than Pythia-14m on quantity estimation—but these gains are neither large nor consistent. In some cases, smaller models like Pythia-160M outperform their larger counterparts in point prediction metrics such as APE. A similar trend appears within the OLMo family: while the OLMo-2-7B variants perform competitively in calibration metrics (e.g., CRPS), this is not accompanied by uniform improvements in accuracy or absolute error.
These observations suggest that increases in parameter count do not automatically translate to improved forecasting performance. While larger models often benefit from greater representational capacity, the forecasting task—particularly when it involves expressing well-calibrated uncertainty—may place demands that are orthogonal to model size. Specifically, forecasting tasks may require targeted reasoning over time, quantities, or epistemic uncertainty, which are not clearly emphasized during standard pretraining or scaled parameter growth. These findings indicate that effective forecasting may depend more on model behavior during decoding or task-specific adaptation, rather than scale alone.
@@
Table 6 compares base and instruct-tuned variants of OLMo-7B and OLMo-2-7B across Boolean, timeframe, and quantity forecasting tasks. For Boolean questions, OLMo-7B-Instruct achieves higher accuracy (0.65 vs. 0.21) and a lower Brier score (0.38 vs. 0.82), indicating better confidence calibration. In timeframe prediction, the instruct-tuned OLMo-2-7B-Instruct improves uncertainty estimation, with an ADE of 0.87 and CRPS of 0.82, compared to 0.9981 and 0.9970 for the base OLMo-2-7B. For quantity estimation, instruct-tuned models have slightly higher APE but lower MAE and CRPS, suggesting better uncertainty calibration. While these results point to potential benefits of instruction tuning for probabilistic reasoning, the sample size remains limited due to the recency of instruction-tuned models, and the observed gains may not be statistically robust.
@@
Table 7 compares different aggregation methods for deriving the final prediction and confidence estimate from the top 10 outputs of Llama-7B. Bayesian Aggregation achieves the highest accuracy (0.58) and F1 score (0.35), suggesting it is the most effective at identifying correct outcomes. However, Weighted Average yields a significantly lower Brier score (0.29), indicating superior confidence calibration compared to others.
Majority Vote, Highest Confidence, and Logit Mean Probability produce comparable accuracy and F1 scores but have noticeably higher Brier scores, suggesting weaker uncertainty estimation. These results highlight that even when point prediction performance is similar, aggregation methods substantially impact confidence reliability. The challenge remains in developing techniques that optimize both accuracy and calibration simultaneously, emphasizing the importance of uncertainty-aware forecasting. In preliminary experiments, we also explored verbalized confidence, prompting the model to explicitly generate a probability (e.g., “I’m 80% confident…”). However, most models either failed to produce enough meaningful numeric estimates or generated inconsistent responses, particularly on harder questions. This highlights a limitation in current instruction-following capabilities when it comes to quantifying uncertainty directly.
A promising direction for future work is to fine-tune models specifically to produce structured confidence outputs, possibly by supervising against calibrated targets or training with ranking-based objectives. Additionally, integrating task-specific priors (e.g., event resolution frequency, temporal proximity) into aggregation could further improve calibration. Hence, joint modeling of correctness and uncertainty remains a key open challenge for real-world deployment of forecasting systems.
Lack of Correlation Between Forecast Date and Model Performance
One might expect language models to perform better on forecasting questions that are closer to their training cutoff date, since more relevant information may have been available during pretraining. For instance, a model like GPT2, with a cutoff in December 2017, would intuitively have an advantage when forecasting 2018 events compared to events in 2020. However, as shown in Appendix E, we observe no clear improvement for near-cutoff questions. In many cases, model performance remains flat or even inconsistent across time, with no systematic degradation as the temporal distance increases.
This observation suggests that forecasting is meaningfully different from retrieval-based tasks such as fact-checking or question answering. While those tasks often benefit from training data that includes the target facts or related content, forecasting involves reasoning under uncertainty about events that had not yet occurred during training. Models are not explicitly trained to perform such reasoning, and pretraining objectives generally do not reward predictions about future developments. The absence of a correlation between cutoff proximity and performance highlights that forecasting requires capabilities beyond knowledge access, including abstraction, probabilistic reasoning, and temporal inference. These factors make the task more complex and reinforce the need for dedicated methods that go beyond static knowledge retrieval.